What length of a copper wire of cross-sectional area 0.01 mm^2 will be needed to prepare a resistance of 1 kΩ? Resistivity of copper = 1.7 × 10^-8 Ω-m.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. f base cu = 1.7 * 10^–8 Ω-m A = 0.01 mm^2 = 0.01 * 10^–6 m2 R = 1 KΩ = 10^3 Ω R = fℓ/a ⇒ 10^3 = 1.7 * 10^-8 * ℓ/10^-6 ⇒ ℓ = 10^3/1.7 = 5.08 * 10^3 m = 0.6 km.